# Disentanglement with Factor Quantized Variational Autoencoders

### v2

@@ 1 Introduction
A recent work proposes QLAE [11], a VAE based model that uses discrete representations as a better inductive bias for disentanglement. Discrete representation learning has gained further importance with the tremendous success of vector quantization for representation learning with VQVAE [12], and been used in recent large generative models like Latent Diffusion Models [13] and DALL-E [14]. While the discrete representations are better suited for expressing categories that shape the observation space [12], Hsu et al. [11] use this rationale for disentanglement, and quantize the continuous representations with a small set of scalar values in order to force the model to assign constant meaning to each value. For each dimension of the representation, Hsu et al. [11] define a separate set of scalar values, referred to as a codebook, and quantize each value using a scalar from its corresponding codebook. The quantizing value is selected based on its proximity to the corresponding variable in the continuous representation. Although Hsu et al. [11] demonstrate that using individual codebooks for each dimension, rather than a single global codebook, improves disentanglement, we identify certain challenges with this approach.
[…]
In this work, we design a novel framework that uses discrete representation learning for disentanglement with a single, global codebook. Different than QLAE, we incorporate further inductive biases to aid disentanglement, and introduce a "total correlation" term to the optimization. As detailed in Section 2, various regularizers have been proposed to enhance disentanglement, with total correlation being one such regularizer. This term modifies the original loss function by adding a penalty that encourages disentangled representations, thereby improving the overall disentanglement of the model [15]. We summarize our contributions as follows:
@@
    We introduce FactorQVAE that originally combines factorization as a regularizer and the discrete VAEs for disentanglement.
@@
    We report the effects of factorization and discretization on disentanglement individually and demonstrate the effectiveness of their novel combination for improved disentangled representations.
@@
    We redesign the training frameworks of two discrete VAE models, VQVAE [12] and dVAE [14], to enhance disentanglement and evaluate their performance.
@@ 2 Related Work
Disentangled Representation Learning: Methods for disentangled representation learning can be categorized by their model types, supervision levels, and independence assumptions [16]. We choose the VAE as the base model architecture over alternatives such as generative adversarial networks (GAN) [17] and diffusion model [18]. Several VAE-based approaches have been proposed for disentanglement, incorporating various assumptions and regularizers. For example, Higgins et al. [19] introduce a penalty coefficient, $\beta$, to the loss component in VAE that affects disentanglement; Kim and Mnih [15] include a total correlation term to enforce disentanglement during optimization; Chen et al. [10] decompose the VAE loss component affecting disentanglement into multiple parts and assign different coefficients to these parts, rather than using a single $\beta$ coefficient, in order to improve disentanglement without sacrificing reconstruction quality; Whittington et al. [9] add biologically inspired activity constraints during training; and Hsu et al. [11] utilize discrete representations to aid in disentanglement.
@@
In our work, we propose an unsupervised VAE-based method that assumes independent generative factors and innovatively integrates discrete representation learning with regularized optimization to enhance disentanglement. As we explicitly address the disentangled representation learning problem which can be used in various applications, generative models such as Latent Diffusion Models [13] that implicitly learn disentangled latent representations with the help of multi-modal inputs for high-quality image generation are not in the scope of our work.
@@ 3.3.2 InfoMEC
In our work, we follow the literature and set $d>F$. Specifically, we set $d=2*F$ where $F$ changes based on the dataset. As a consequence of this inequality, achieving both perfect disentanglement (modularity) and perfect completeness (compactness) is impossible. Therefore, achieving a better disentanglement (modularity) value is prioritized.
@@ 4 Method
We hypothesize that using a lower dimensional latent representation $z_{e}(x)\in\mathbb{R}^{d\times 1}$ and quantizing scalar latent variables, rather than using a higher dimensional latent representation $z_{e}(x)\in\mathbb{R}^{d*C\times 1}$ and quantizing embeddings in the latent, can better maintain a balance between reconstruction and disentanglement performance. This is because a higher dimensional latent representation, quantized with the embeddings from a codebook, has a higher information capacity. This increased capacity may shift the emphasis towards improving reconstruction performance during training, potentially compromising disentanglement performance. Therefore, we proceed with scalar quantization in the proposed method.
@@
Since the closest scalar in $\mathcal{M}$ has the lowest distance, we use $-\mathcal{R}$ as the parameters of the Categorical distribution to be able to sample the closest scalar with a higher probability in Equation 4. Sampling from the stochastic posterior $q(z|x)$ yields the discrete variables $z\in\mathbb{R}^{d\times K}$ which are essentially the one-hot representations of the indices of the quantizing scalars for $d$ number of latent variables. Since sampling from a Categorical distribution is not a differentiable operation in the traditional sense that hurts end-to-end training, we use Gumbel-Softmax distribution for a differentiable approximation to categorical sampling [33, 34]. These differentiable samples will be essential when we add the total correlation term to the optimization.
@@
     $\mathcal{M}^{[t]},\theta_{1}^{[t]},\theta_{2}^{[t]},\phi_{1}^{[t]},\phi_{2}^{[t]}\leftarrow g_{1}(\nabla_{\mathcal{M},\theta_{1},\theta_{2},\phi_{1},\phi_{2}}\mathcal{L}_{1},\text{learning rate}=\alpha^{[t-1]}_{1})$
@@
     $\psi^{[t]}\leftarrow g_{2}(\nabla_{\psi}\mathcal{L}_{2},\text{learning rate}=\alpha^{[t-1]}_{2})$
@@ 5.1 Experimental Settings
Table 1: Evaluation of models across all datasets for disentanglement (DCI and InfoMEC) and reconstruction (MSE ×$\times 10^{4}$) performance.
@@
We conduct our experiments on three datasets: Shapes3D [35], Isaac3D [36], and MPI3D [37]. Each dataset corresponds to simple, medium, hard complexity, respectively. Further details about the datasets are given in Appendix A.1.
We choose recent and state-of-the-art disentanglement models to compare our model: $\beta$-VAE [19], FactorVAE [15], BioAE [9], and QLAE [11]. We also evaluate autoencoder (AE), QVAE which is the scalar quantizing version of VQVAE [12] instead of vector quantization, and dVAE [14]. We use the same network architecture across all baseline models to ensure a fair comparison. While this may result in minor deviations from the original findings reported in the literature, these differences are not significant enough to impact the overall conclusions. Further details like the network architecture and hyperparameter selection for each model are given in Appendix A.3.
@@ 5.2 Discussions
Figure 2: Latent traversal on the same image with QLAE and FactorQVAE for Shapes3D dataset. Each row $i$ shows the result of manipulating the $i^{th}$ latent. For QLAE, the $i^{th}$ latent is intervened on with a linear interpolation between the minimum and maximum values in the corresponding $i^{th}$ codebook while it is intervened on with a linear interpolation between the minimum and maximum values in the global codebook for FactorQVAE.
@@
Figure 3: Latent traversal on the same image with FactorVAE and FactorQVAE for Isaac3D dataset. Each row $i$ shows the result of manipulating the $i^{th}$ latent. For FactorVAE, the $i^{th}$ latent is intervened on with a linear interpolation between "original latent value - 3" and "original latent value + 3" while it is intervened on with a linear interpolation between the minimum and maximum values in the global codebook for FactorQVAE.
@@
Figure 4: Latent traversal on the same image with $\beta$-VAE and FactorQVAE for MPI3D dataset. Each row $i$ shows the result of manipulating the $i^{th}$ latent. For $\beta$-VAE, the $i^{th}$ latent is intervened on with a linear interpolation between "original latent value - 3" and "original latent value + 3" while it is intervened on with a linear interpolation between the minimum and maximum values in the global codebook for FactorQVAE.
@@
To evaluate the disentanglement numerically, we use DCI and InfoMEC, with details provided in Section 3.3. Table 5.1 shows both DCI and InfoMEC results of the models for all datasets. As disentanglement should be obtained along with acceptable reconstruction performance, we further present the mean squared error (MSE) of the models in Table 5.1. To highlight our key results, we draw attention to the $\mathsf{D}$ and InfoM scores in Table 5.1, where our method clearly outperforms the baselines in most of the cases. While $\mathsf{C}$ and InfoC scores may not show a similar level of improvement, this aligns with our discussion in Section 3.3, where we prioritize disentanglement (modularity) to completeness (compactness) due to the inherent trade-off dictated by the chosen parameter settings ($d=2*F$).
@@
Discrete vs continuous representation learning: To start with, we compare discrete representation learning and continuous representation learning in terms of disentanglement and reconstruction. When we compare models that learn discrete representations, such as QLAE and QVAE, with models that learn continuous representations, such as $\beta$-VAE and FactorVAE, by referring to Table 5.1, we observe that while the discrete representation learning models outperform the continuous representation learning models in terms of the DCI metric for all datasets, they also outperform the latter in terms of InfoMEC in most cases. In addition to disentanglement performance, the discrete representation learning models achieve better reconstruction performance. Since achieving both acceptable reconstruction and disentanglement performance together is difficult, an AE trained solely on the reconstruction objective provides the best reconstruction performance, as expected. However, discrete representation learning models can also achieve reconstruction performance that is acceptably close to that of autoencoders (AE).
@@
Scalar quantization vs vector quantization: In order to test our intuition behind scalar quantization explained in Section 4, we form various codebook configurations, and test them using Shapes3D dataset. Table 5.2 presents the results of our experiments. Firstly, we focus on the effects of scalar quantization. We name the models VQVAE and FactorVQVAE when the codebook $\mathcal{M}$ consists of embeddings, and we perform vector quantization. We observe that going from scalars to embeddings, and further increasing the dimensionality of the embeddings leads to a slightly better reconstruction performance as expected. However, we monitor a notable decrease in disentanglement, especially in terms of InfoMEC. Therefore, we experimentally confirm that scalar quantization leads to better maintain a balance between the reconstruction and disentanglement performance. Moreover, we experience a stability problem when we train VQVAE and FactorVQVAE with $\mathcal{M}\in\mathbb{R}^{64\times 8}$ and $\mathcal{M}\in\mathbb{R}^{64\times 8}$ as the training is more vulnerable to changes in $\beta$ and $\gamma$ coefficients. Due to its demonstrated benefits, we opted for scalar quantization.
@@ 6 Conclusion
The proposed FactorQVAE extends the discrete representation learning model QVAE by introducing a regularizer called factorization to enhance disentanglement in representation learning. First, we show that discrete representation learning is more suitable than continuous representation learning for disentanglement. We further demonstrate that incorporating an inductive bias into a discrete representation learning model improves its performance, with FactorQVAE outperforming previous methods in most settings in terms of both disentanglement and reconstruction.

### v3

@@ 1 Introduction
A recent work proposes QLAE [11], a VAE based model that uses discrete representations as a better inductive bias for disentanglement. Discrete representation learning has gained further importance with the tremendous success of vector quantization for representation learning with VQVAE [12], and has been used in recent large generative models like Latent Diffusion Models [13] and DALL-E [14]. While discrete representations are better suited for expressing categories that shape the observation space [12], Hsu et al. [11] use this rationale for disentanglement, and quantize the continuous representations with a small set of scalar values in order to force the model to assign constant meaning to each value. For each dimension of the representation, Hsu et al. [11] define a separate set of scalar values, referred to as a codebook, and quantize each value using a scalar from its corresponding codebook. The quantizing value is selected based on its proximity to the corresponding variable in the continuous representation. Although Hsu et al. [11] demonstrate that using individual codebooks for each dimension, rather than a single global codebook, improves disentanglement, we identify certain challenges with this approach.
In QLAE, the quantization of each latent variable is restricted to its own codebook that includes a predefined number of scalar values, k. Assume two generative factors, shape and orientation that are ideally captured with different latent variables. These factors have a different number of options, n and m where $\textbf{n}<\textbf{m}$, respectively. Therefore, the range of values in the latent representation may vary to capture these options, such that the range of latent values for orientation might be wider since $\textbf{n}<\textbf{m}$. Quantizing the latent values of different generative factors using individual codebooks can be limiting in terms of representation capacity. For example, using k scalars may be insufficient to represent m possible orientations, while at the same time, k could be excessive for representing n possible shapes. Even if the codebook size k is large enough to cover the generative factor with the most options, this would lead to redundancy in the codebook for other generative factors with fewer options, as the codebook size remains equal for each factor.
Although QLAE’s use of individual codebooks promotes disentanglement, its fixed-size, per-dimension quantization introduces inefficiencies and limits representational flexibility, especially when generative factors have diverse cardinalities. To overcome these limitations, we propose FactorQVAE, a novel discrete latent variable framework explicitly designed for disentangled representation learning. Unlike prior discrete models such as VQVAE and dVAE, which primarily enhance reconstruction quality or sampling efficiency, FactorQVAE integrates scalar quantization using a single global codebook with a total correlation regularizer. As discussed in Section 2, total correlation is a prominent regularizer for disentanglement, encouraging statistical independence among latent variables by adding a specific penalty term to the loss function [15]. Integrating scalar quantization with total correlation is challenging because discrete latents must remain differentiable and limited codebook capacity must be managed, which FactorQVAE addresses via a stochastic categorical posterior and a global codebook. These innovations translate directly into practical benefits for real-world tasks like robotics, visual reasoning, and fairness-aware systems. For instance, disentangling object attributes facilitates robotic adaptability, while isolating sensitive factors enhances transparency in fairness-critical systems. Thus, FactorQVAE bridges fundamental theoretical advances with meaningful practical applications.
We summarize our contributions as follows:
@@
    We introduce FactorQVAE that originally combines factorization as a regularizer and the discrete VAEs for disentanglement, using scalar quantization with a single global codebook to overcome the representational inefficiencies of fixed per-dimension codebooks.
@@
    We report the effects of factorization and discretization on disentanglement individually and demonstrate the effectiveness of their novel combination through a stochastic categorical posterior for improved disentangled representations.
@@
    We redesign the training frameworks of two discrete VAE models, VQVAE [12] and dVAE [14], to enhance disentanglement and systematically evaluate the trade-offs between scalar and vector quantization approaches.
@@ 2 Related Work
Disentangled Representation Learning: Methods for disentangled representation learning can be categorized by their model types, supervision levels, and independence assumptions [16]. We choose the VAE as the base model architecture over alternatives such as generative adversarial networks (GAN) [17] and diffusion models [18].
Several VAE-based disentanglement approaches have been proposed, each with limitations motivating our work. Higgins et al. [19] introduce $\beta$-VAE, where a higher $\beta$ creates a fundamental reconstruction–disentanglement trade-off, significantly reducing reconstruction fidelity and producing uninformative latent representations. Kim and Mnih [15]’s FactorVAE partially mitigates these issues by explicitly penalizing total correlation, yet the hyperparameter-sensitive trade-off persists. Chen et al. [10] propose decomposing the VAE loss with multiple separate hyperparameters, which enhances flexibility but exacerbates hyperparameter tuning challenges. Whittington et al. [9]’s BioAE incorporates biologically inspired constraints, which introduce domain-specific assumptions limiting its generalizability. QLAE proposed by Hsu et al. [11], another discrete latent representation approach, requires unusually strong regularization (e.g., high weight decay), and its per-dimension codebooks restrict latent expressiveness.
Recent methods also face distinct practical challenges. Meo et al. [20]’s $\alpha$-TCVAE combines variational and conditional entropy bottlenecks but remains sensitive to the $\alpha$ hyperparameter balancing diversity and reconstruction quality. Li et al. [21]’s PDisVAE leverages partial correlation for group-wise independence but struggles with complex hyperparameter optimization and selecting optimal latent groupings. Uscidda et al. [22]’s approach based on the Gromov-Monge Gap regularizer using optimal transport theory suffers from quadratic computational complexity and demands careful hyperparameter tuning.
@@
In contrast, our proposed FactorQVAE is an unsupervised VAE-based method explicitly assuming independent generative factors and uniquely integrating discrete latent quantization with total correlation regularization. Unlike prior methods, FactorQVAE simultaneously optimizes stability, reconstruction quality, and disentanglement, significantly mitigating hyperparameter sensitivity. This translates directly into practical advantages, including more robust performance across diverse datasets and easier training procedures.
Note that generative models such as Latent Diffusion Models [13], which implicitly leverage disentanglement through multimodal inputs for high-quality image generation, fall outside our scope, as our focus remains explicitly on structured disentangled representations. Likewise, recent approaches [26, 27] that prioritize disentanglement for only selected latent dimensions while allowing others to remain entangled address a different aspect of the problem. In contrast, FactorQVAE relies on discrete quantization and treats all latent dimensions uniformly, ensuring balanced disentanglement across the representation. In addition, recent work has explored disentanglement in broader domains [28, 29], reflecting a growing interest in adapting disentangled representation learning beyond 3D-structured benchmarks. Our additions position FactorQVAE within this broader context while highlighting its distinct contribution in discrete, factorized representation learning.
@@ 3.3.2 InfoMEC
In our work, we follow the literature and set $d>F$. Specifically, we set d=2Fd=2*F where $F$ changes based on the dataset. As a consequence of this inequality, achieving both perfect disentanglement (modularity) and perfect completeness (compactness) is impossible. Therefore, achieving a better disentanglement (modularity) value is prioritized.
@@ 4 Method
In designing our framework, we intentionally employ scalar quantization instead of alternative approaches such as vector quantization. Scalar quantization explicitly constrains the latent representation’s information capacity, promoting an effective balance between reconstruction and disentanglement. In contrast, vector quantization provides higher-dimensional latent embeddings with greater representational capacity, which can inadvertently lead the model to prioritize reconstruction accuracy over disentangled representations by encoding redundant or correlated information. From an information-theoretic perspective, this difference can be understood as follows: vector quantization allows each codebook vector to capture correlations across multiple latent dimensions, thereby increasing inter-dimensional mutual information and making disentanglement more difficult. Scalar quantization, by contrast, enforces independent per-dimension assignments, which act as a stronger information bottleneck. This encourages alignment between individual latent variables and generative factors while reducing redundancy between dimensions. Hence, our deliberate choice of scalar quantization provides both an intuitive and theoretically motivated inductive bias toward disentanglement. We validate this decision empirically in Section 5 and Table 5.2.
@@
Since the closest scalar in $\mathcal{M}$ has the lowest distance, we use -$\mathcal{R}$ as the parameters of the Categorical distribution to be able to sample the closest scalar with a higher probability in Equation 4. Sampling from the stochastic posterior $q(z|x)$ yields the discrete variables $z\in\mathbb{R}^{d\times K}$ which are essentially the one-hot representations of the indices of the quantizing scalars for $d$ number of latent variables. Sampling from a Categorical distribution is inherently non-differentiable, which complicates end-to-end training when combined with the total correlation regularizer. To overcome this, we adopt the Gumbel–Softmax distribution as a differentiable approximation [40, 41], allowing gradient-based optimization while preserving the discrete nature of the latent variables. These differentiable samples will be essential when we add the total correlation term to the optimization.
@@
[…]
@@
[…]
@@ 5.1 Experimental Settings
Table 1: Evaluation of models across all datasets for disentanglement (DCI and InfoMEC) and reconstruction (MSE $\times 10^{4}$) performance.
@@
We conduct our experiments on three datasets: Shapes3D [42], Isaac3D [43], and MPI3D [44]. Each dataset corresponds to simple, medium, hard complexity, respectively. In our experiments, MPI3D refers to the MPI3D-Complex variant, which contains images captured from a real robotic arm and physical objects under varying backgrounds and lighting conditions, making it representative of real-world scenarios. Further details about the datasets are given in Appendix A.1.
We choose fundamental, recent, and state-of-the-art disentanglement models to compare our model: $\beta$-VAE [19], FactorVAE [15], BioAE [9], $\alpha$-TCVAE [20], and QLAE [11]. We also evaluate autoencoder (AE), QVAE which is the scalar quantizing version of VQVAE [12] instead of vector quantization, and dVAE [14]. We use the same network architecture across all baseline models to ensure a fair comparison. While this may result in minor deviations from the original findings reported in the literature, these differences are not significant enough to impact the overall conclusions. Further details, such as hyperparameter selection and ablations, are provided in Appendix A.2, and the network architecture is described in Appendix A.3.
@@ 5.2 Discussions
Figure 2: Latent traversal on the same image with QLAE and FactorQVAE for Shapes3D dataset. Each row $i$ (labeled as $\mathbf{z_{i}}$) shows the result of manipulating the $i^{th}$ latent, with the last 8 columns showing interpolations. For QLAE, the $i^{th}$ latent is intervened on with a linear interpolation between the minimum and maximum values in the corresponding $i^{th}$ codebook while it is intervened on with a linear interpolation between the minimum and maximum values in the global codebook for FactorQVAE. For FactorQVAE, rows 1 and 2 control object hue, row 3 controls object shape, rows 4 and 10 control camera orientation, rows 5 and 11 control floor hue, rows 6 and 8 control wall hue, and rows 7 and 9 control object scale.
@@
Figure 3: Latent traversal on the same image with FactorVAE and FactorQVAE for Isaac3D dataset. Each row $i$ (labeled as $\mathbf{z_{i}}$) shows the result of manipulating the $i^{th}$ latent, with the last 8 columns showing interpolations. For FactorVAE, the $i^{th}$ latent is intervened on with a linear interpolation between "original latent value - 3" and "original latent value + 3" while it is intervened on with a linear interpolation between the minimum and maximum values in the global codebook for FactorQVAE. For FactorQVAE, rows 1 and 6 control lighting direction, rows 3 and 10 control the robot’s vertical axis, rows 5 and 8 control object scale, rows 9 and 13 control wall color, rows 10 and 14 control object color, row 11 controls the robot’s horizontal axis, row 12 controls lighting intensity, row 16 controls camera height, and row 17 controls object shape.
@@
Figure 4: Latent traversal on the same image with $\beta$-VAE and FactorQVAE for MPI3D dataset. Each row $i$ (labeled as $\mathbf{z_{i}}$) shows the result of manipulating the $i^{th}$ latent, with the last 8 columns showing interpolations. For $\beta$-VAE, the $i^{th}$ latent is intervened on with a linear interpolation between "original latent value - 3" and "original latent value + 3" while it is intervened on with a linear interpolation between the minimum and maximum values in the global codebook for FactorQVAE. For FactorQVAE, row 2 controls camera height, row 3 controls object color, row 5 controls object size, row 8 controls background color, and row 9 controls object shape.
@@
To evaluate the disentanglement numerically, we use DCI and InfoMEC, with details provided in Section 3.3. Table 5.1 shows both DCI and InfoMEC results of the models for all datasets. As disentanglement should be obtained along with acceptable reconstruction performance, we further present the mean squared error (MSE) of the models in Table 5.1. To highlight our key results, we draw attention to the $\mathsf{D}$ and InfoM scores in Table 5.1, where our method clearly outperforms the baselines in most of the cases. While $\mathsf{C}$ and InfoC scores may not show a similar level of improvement, this aligns with our discussion in Section 3.3, where we prioritize disentanglement (modularity) to completeness (compactness) due to the inherent trade-off dictated by the chosen parameter settings (d=2Fd=2*F).
@@
Discrete vs continuous representation learning: To start with, we compare discrete representation learning and continuous representation learning in terms of disentanglement and reconstruction. By comparing discrete representation models (FactorQVAE, QLAE, and QVAE) against continuous representation methods ($\beta$-VAE, FactorVAE, and $\alpha$-TCVAE) in Table 5.1, we observe that discrete latent models consistently outperform continuous methods in both DCI and InfoMEC metrics across datasets. In addition to disentanglement performance, the discrete representation learning models achieve better reconstruction performance. Since achieving both acceptable reconstruction and disentanglement performance together is difficult, an AE trained solely on the reconstruction objective provides the best reconstruction performance, as expected. However, discrete representation learning models can also achieve reconstruction performance that is acceptably close to that of autoencoders (AE).
@@
Scalar quantization vs vector quantization: To substantiate the methodological choice detailed in Section 4, we conduct an ablation study comparing scalar and vector quantization strategies using the Shapes3D dataset. Results in Table 5.2 demonstrate that vector quantization marginally improves reconstruction but substantially reduces disentanglement scores (e.g., InfoMEC metric). Additionally, we observe increased training instability in vector quantized models (e.g., VQVAE and FactorVQVAE with $\mathcal{M}\in\mathbb{R}^{64\times 8}$), which are more sensitive to hyperparameter variations. These findings confirm our hypothesis that scalar quantization provides a more suitable trade-off, effectively balancing disentanglement and reconstruction performance.
@@ 6 Conclusion
The proposed FactorQVAE extends the discrete representation learning model QVAE by introducing a regularizer called factorization to enhance disentanglement in representation learning. First, we show that discrete representation learning is more suitable than continuous representation learning for disentanglement. We further demonstrate that incorporating an inductive bias into a discrete representation learning model improves its performance, with FactorQVAE outperforming previous methods in most settings in terms of both disentanglement and reconstruction. Since our evaluation focuses on standard 3D-object benchmarks in disentanglement research, the conclusions are primarily applicable to structured visual domains with well-defined generative factors. 3D scenes are a particularly suitable testbed since their generative factors are naturally interpretable and controllable, making them central to disentanglement research. Extending FactorQVAE to less structured datasets or non-visual modalities remains a promising direction for future work.